how do i solve this equation 3x+5=3-2x?
Found 2 solutions by checkley77, mickclns:
Answer by checkley77(12844)

3x+5=3-2x
3x+2x=3-5
5x=-2
x=-2/5 ans.
Proof:
3(-2/5)+5=3-2(-2/5)
-6/5+5=3+4/5
(-6+25)/5=(3*5+4)/5
19/5=(15+4)/5
19/5=19/5


Answer by mickclns(59)

3x+5=3-2x add 2x to both sides to make sure all x's are on the same side.
5x + 5 = 3 subtract 5 from both sides to make sure all the constants are NOT on the same side as x.
5x = -2 to isolate x, divide both sides by 5
x = -2/5